OpenStudy (anonymous):

What is 1550 dm expressed in hectometers?

OpenStudy (johnweldon1993):

There are 1000 dm in 1 hm so we have \[\large 1550 dm \times \frac{1 hm}{1000 dm} = \frac{1550}{1000} = 1.55hm\]